How to calculate gram equivalent weight of potassium dichromate ?

Grade:12

1 Answers

Umakant biswal
7 years ago
dear jayasree 
molecular weight of pottasium dichromate ( k2cr2o7 ) = 294 g 
X factor in acidic medium = 6 ( as there is nothing mentioned in ques i assumed it to be acidic ) 
equivalent weight = molecular weight / X factor 
= 294 / 6 = 49 g 
thats your answer 
HOPE IT CLEARS 
REGARDS